Hi, Can I please have some help with this problem about functions? thank you!
An appliance dealer advertises a 15% discount on all his washing machines. In addition, the manufacturer offers a $80 rebate on the purchase of a washing machine. Let x represent the sticker price of the washing machine.
(a) Suppose only the 15% discount applies. Fund a function f that models the purchase price of the washer as a function of the sticker price x
f(x) =
(b) Suppose only the $80 rebate applies. Find a function g that models the purchase price of the washer as a function of the sticker price x.
g(x) =
(c) Find f º g.

(d) Find g º f.

Answer by stanbon(75887)

An appliance dealer advertises a 15% discount on all his washing machines. In addition, the manufacturer offers a $80 rebate on the purchase of a washing machine. Let x represent the sticker price of the washing machine.
(a) Suppose only the 15% discount applies. Fund a function f that models the purchase price of the washer as a function of the sticker price x
f(x) = 0.85x
------------------------
(b) Suppose only the $80 rebate applies. Find a function g that models the purchase price of the washer as a function of the sticker price x.
g(x) = x - 80
-------------------
(c) Find f º g(x) = f[x-80] = 0.85(x-80) = 0.85x-0.68
------------------
(d) Find g º f(x) = g[0.85x] = 0.85x-80
-----------------------------------
Cheers,
Stan H.